The thick lens shown at right is made from glass with index of refraction 1.50. The thick lens has thickness of 6.0 cm and radii of curvature of 8.00 cm and 10.00 cm. (a) Find the focal length of this thick lens in air and the position of its focal points (F and F) and its principal points (H, and H,) with respect to the front and back of the lens respectively. (b) Draw a sketch showing the locations of the focal points and the principle planes. Indicate all distance on the sketch. 1.
Answer #1

thicknes (d)6 0:5 (D.125-0.1十0.02.5 = pr.nci peialt») =-I(ne-l--D·D25G 5DE mR 2 mean f0.02 cm -Back focal PonH 0.025= 2 +6.25thick lens MiN Prinal plans